OpenStudy (anonymous):

Another LaTeX doc question: How do say this?... Ok, In a Word document I would just hit the enter key and I can write on a new line. In latex, on the other had, it doesn't start a new line for me.

hero (hero):

In latex, you start a new line by writing this `\[\text{1st Line}\]\[\text{2nd Line}\]` It will look like this afterwards: \[\text{1st Line}\]\[\text{2nd Line}\]

hartnn (hartnn):

new line in LATEX `use \\` example `\( a= b+c \\ a-b=c\)` gives \( a= b+c \\ a-b=c\)

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

How about space between paragraphs?

hero (hero):

`\space`

hero (hero):

Matter of fact, maybe you can use @hartnn's methods to do that: `\[\text{1st Line} \\ \space \\ \text{3rd Line}\]` \[\text{1st Line} \\ \space \\ \text{3rd Line}\]

OpenStudy (anonymous):

oh I see. Cool, thanks!